ABC Corporation manufactures lamps. It has set up the following standards per finished unit for direct materials and direct manufacturing labor:

Direct materials: 2 lb. at $7.50 per lb.

$15.00

Direct manufacturing labor: 0.3 hour at $90 per hour

$27.00

The number of finished units budgeted for January 2017 was 20,000; 15000 units were actually produced.

Actual results in January 2017 were as follows:

Direct materials used: 2.2 lb. x 15,000 = 33,000 lbs @ $7.00 / lb.

Direct manufacturing labor: 0.32 hrs x 15,000 = 4,800 hrs @ $95 / hr

Assume that there was no beginning inventory of either direct materials or finished units.

The following data are for January 2017:

Budgeted

Actual

Number of Lamp sold

20,000 units

15,000

Selling price / unit

$80

$85

Variable cost per unit

Direct material

(2 lbs x $7.50 = $15)

(2.2 lbs x $7 = $15.40)

Direct labor

(0.3 hrs x $90 = $27)

(0.32 hrs x $95 = $30.40)

Fixed cost

$240,000

$240,000

During the month, materials purchased amounted to 40,000 lb. @ $7.00, at a total cost of $280,000. Input price variances are isolated upon purchase. Input-efficiency variances are isolated at the time of usage.

Required:

1.   Compute the Static-Budget variance, Flexible-Budget variance, and Sales-volume variance for January 2018 in the table below:

2.   Compute the price and efficiency variances of direct material and direct manufacturing labor for January 2017 in the table below.

Answer

1. DMPV

2. DMEV

3. DLRV

4. DLEV

3.   Please provide 3 possible reasons for the material price variance and material efficiency variance.

Actual

Results

(1)

Flexible-

Budget

Variances

(2) = (1) – (3)

Flexible

Budget

(3)

Sales

Volume

Variances

(4) = (3) – (5)

Static

Budget

(5)

Static

Budget

Variance

Units sold

Revenue

Variable costs

DM

DL

Total variable costs

Contribution margin

Fixed costs

Operating income
Answer #1
1 Budget Variances
Actual Results Flexible Budget Variance Flexible Budget Sales Volume Variance Static Budget Static Budget Variance
.(1) .(2)=.(1)-(.3) .(3) .(4)=(.3)-.(5) .(5) .(6)=.(1)-.(5)
a Units Sold 15000 15000 20000
b Selling Price $85 $80 $80
c=a*c Revenue $1,275,000 $75,000 $1,200,000 ($400,000) $1,600,000 ($325,000)
Variable Costs:
d Direct material per unit $15.40 $15.00 $15.00
e=a*d Total Direct Materials Costs $231,000 $6,000 $225,000 ($75,000) $300,000 ($69,000)
f Direct labor per unit $30.40 $27.00 $27.00
g=a*f TotalDirect Labor Cost $456,000 $51,000 $405,000 ($135,000) $540,000 ($84,000)
h=e+g Total Variable Cost $687,000 $57,000 $630,000 ($210,000) $840,000 ($153,000)
i=c-h Contribution margin $588,000 $18,000 $570,000 ($190,000) $760,000 ($172,000)
j Fixed Costs $240,000 $0 $240,000 $0 $240,000 $0
k=i-j Operating Income $348,000 $18,000 $330,000 ($190,000) $520,000 ($172,000)
2 Material and Labor Variances
DIRECT MATERIALS VARIANCES
SQ Standard Quantity(Pounds) for actual output 30000 (2*15000)
SP Standard Price per Pound $7.50
AQ Actual Quantity of materials used(Pounds) 33000 (2.2*15000)
AP Actual Price per unit paid $7.00 (121000/10000)
DMPV AQ*(AP-SP) Direct Materials Price Variance $16,500 Favorable (Actual price is lower than Standard Price)
DMEV SP*(AQ-SQ) Direct Materials Quantity Variance $22,500 Un Favorable (Actual quantity is more than Standard quantity)
DIRECT LABOR VARIANCES
SH Standard Labor hour for actual output 4500 (0.3*15000)
SR Standard Labor rate per hour $90.00
AH Actual labor hour used 4800 (0.32*15000)
AR Actual Labor rate per hour $95.00
DLRV AH*(AR-SR) Direct labor Rate Variance $24,000 Unfavorable (Actual Rate is higher than Standard Rate)
DLEV SR*(AH-SH) Direct Labor Efficiency   Variance $27,000 Un favorable (Actual hour is more than Standard hour)